OpenStudy (anonymous):

Find the value of x so that the line passing through (x, 10) and (-4, 8) has a slope of 2/3.

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) 2/3 = (8 - 10)/(-4 - x) 2/3 = (-2)/(-4 - x) 2(-4 - x) = 3(-2) -8 - 2x = -6 I'll let you finish